# dense matter, dark matter, compact star

### v1

@@ 3 Strangeon stars
There are a lot of mass-radius observations to support the SnS model. The photospheric radius expansion bursts of 4U 1746-37 constrain the mass-radius range of low-mass pulsars, and strangeon-star equation of state (Lai and Xu, 2009) meets the constraint better than neutron-star equations of state (Li et al., 2015), see 3. The discovery of massive neutron stars with mass $M\gtrsim 2M_{\odot}$ such as PSR J1614-2230 (Demorest et al., 2010), PSR J0740+6620 (Cromartie et al., 2020), and PSR J0952-0607 (Romani et al., 2022) supports a stiffer equation of state. As seen in Figure 3, SnM is stiffer and can reach a $3M_{\odot}$ maximum mass.
@@
Figure 3: The mass-radius curve calculated using Tolman-Oppenheimer-Volkoff equation and equations of state, This figure is from (Lai et al., 2025). The first five types in the legend is the different kinds of equations of state of SnSs. The MIT bag model is a quark state, and the AP4 is a neutron-proton-electron-muon state.
Furthermore, the SnM state can help us to understand subpulses in some single pulses. Because the base unit of SnM is non-relativistic strangeons, the strong interaction between strangeons is strong enough to make the matter solid. This solid SnM is different from a neutron star’s solid crust because the latter is mainly combined by electromagnetic interaction. So, an SnS could support a higher and longer lasting “zit” (little hill) than a neutron star’s crust could support. These zits may be the spark origin of some highly linear-polarized subpulses (Vivekanand and Radhakrishnan, 1981; Cao et al., 2024; Wang et al., 2024), see Figure 4.
@@ 4 Strangeon nuggets
The SnM can also form small nuggets, called strangeon nuggets (SnNs). The baryon number of smallest stable SnN, i.e. the critical baryon number $A_{\mathrm{c}}$, is uncertain due to the difficulty of computation of the non-perturbative strong interaction. If one considers the role of the electron in breaking the symmetry of $e^{\pm}$ (Xu, 2020), the critical baryon number can be estimated from the Compton wavelength of the electron, $A_{\mathrm{c}}\approx(\lambda_{\mathrm{C}}/r_{\rm{nucleon}})^{3}\approx 10^{10}$. We have only ever observed strong interacting matter with baryon number $A<300$ (atomic nuclei) or $A\approx 10^{57}$ (pulsars). If this assumption holds, the mass spectrum of strong interacting matter can be broadened considerably (Wu et al., 2024; Xu, 2024), see Figure 5.
@@
SnNs with $10^{25}<A<10^{35}$, as quark nuggets, are among the candidates for dark matter. Since the light-three-flavor symmetry in SnM is only slightly broken by the larger strange-quark mass, the charge-to-mass ratio of SnM is low, the electromagnetic signals from SnN are too weak to be observed. SnNs move virially in the universe at a typical speed of about $10^{-3}c$, when they hit Earth, they may not generate a high-energy shower signal, but rather an acoustic shock wave. So, we can detect these non-relativistic nuggets using the acoustic method (Qi and Xu, 2025). Due to the extremely high density of SnNs, their penetration depth in rock or water is very large, even exceeding the radius of the Earth, the calculation results is shown in Figure 6, in which the dimensionless $k$ factor represents the ratio of the assumed interaction cross-section between the SnN and the ordinary matter and the geometric cross-section of the SnN.
@@
The initial shock wave is strong but will decay during propagation in the medium due to diffusion, dispersion, dissipation and scattering. The decay can be described as an exponential function. At a critical distance $R_{\rm crit}$, the acoustic signal decay to the level of surrounding noise, $R_{\rm crit}$ is the maximum detectable distance. This distance depends on the state, the temperature and the uniformity of the medium. An estimated result is in Figure 7, $R_{\rm crit}\simeq 30\,$km. More accurate results require on-site acoustic experiments.

### v2

@@ 3 Strangeon stars
There are a lot of mass-radius observations to support the SnS model. The photospheric radius expansion bursts of 4U 1746-37 constrain the mass-radius range of low-mass pulsars, and strangeon-star equation of state (Lai and Xu, 2009) meets the constraint better than neutron-star equations of state (Li et al., 2015), see Figure 3. The discovery of massive neutron stars with mass $M\gtrsim 2M_{\odot}$ such as PSR J1614-2230 (Demorest et al., 2010), PSR J0740+6620 (Cromartie et al., 2020), and PSR J0952-0607 (Romani et al., 2022) supports a stiffer equation of state. As seen in Figure 3, SnM is stiffer and can reach a $3M_{\odot}$ maximum mass.
@@
Figure 3: The mass-radius curve calculated using Tolman-Oppenheimer-Volkoff equation and equations of state, the theoretical limits from general relativity, and the observational constraints from 4U 1746-37. This figure is from (Li et al., 2015). The green lines and red lines are from the strangeon-star equations of state with different parameters, the purple dotted-dashed line is from the MIT bag equation of state, the black curves are from the neutron-star equations of state. The black dashed line is $M=2M_{\odot}$, due to the presence of PSR J0740+6620, the highest point of the MR curve should be higher than this line. The gray regions from deep to shallow are $1\sigma$, $2\sigma$, $3\sigma$ mass-radius confidence regions of 4U 1746-37 based on the assumption that the touchdown flux and emission area were partially obscured by the accretion disk. It can be seen that the MIT bag quark star and strangeon star models better conform to 4U 1746-37 observation than neutron stars, while the strangeon star model has a larger maximum mass and is more suitable for observing large-mass neutron stars like PSR J0740+6620.
Furthermore, the SnM state can help us to understand subpulses (referring to a short signal structure within a single pulse cycle) in the observed single pulse signals. Because the base unit of SnM are non-relativistic strangeons, the strong interaction between strangeons is strong enough to make the matter solid at low temperature. Due to the large rest mass of the strangeon, its de Broglie wavelength is relatively small, so it behaves more like a classical particle. Taking an example of electromagnetic matter, when inert gases are cooled at standard atmospheric pressure, only helium will exhibit superfluidity, and heavier elements will combine to form classical solids. Assuming that the Lennard-Jones potential is used to describe the interaction between strangeons, Bayesian analysis of observations shows that the potential well depth of the interaction potential between adjacent strangeons can reach over 10 MeV (Yuan et al., 2025). When the temperature of strangeon matter is much lower than 10 MeV (which is met by most of neutron stars), we believe that strangeon matter will also combine into classical solids. This solid SnM is stiffer than a neutron star’s solid crust because the former is combined by strong interaction, but the latter is mainly combined by electromagnetic interaction. So, an SnS could support a higher and longer lasting “zit” (little hill) than a neutron star’s crust could support. These zits may be the spark origin of some highly linear-polarized subpulses (Vivekanand and Radhakrishnan, 1981; Cao et al., 2024; Wang et al., 2024), see Figure 4.
@@ 4 Strangeon nuggets
The SnM can also form small nuggets, called strangeon nuggets (SnNs). The baryon number of smallest stable SnN, i.e. the critical baryon number $A_{\mathrm{c}}$, is uncertain due to the difficulty of computation of the non-perturbative strong interaction. If one considers the role of the electron in breaking the symmetry of $e^{\pm}$ (Xu, 2020), the critical baryon number can be estimated from the Compton wavelength of the electron, $A_{\mathrm{c}}\approx(\lambda_{\mathrm{C}}/r_{\rm{nucleon}})^{3}\approx 10^{10}$. In fact, we have not yet conducted a very detailed calculation on the analysis of weak effects on the stability of strangeon matter, because the strong interaction plays a decisive role in the absolute stability of strange matter in terms of energy. Like the Bodmer-Witten assumption, we assume that the energy of bulk strangeon matter at zero temperature and zero pressure is less than 930 MeV. For little strangeon nuggets with $A<A_{\mathrm{c}}$, protons and neutrons evaporate from them. At that time, finite volume effects such as surface tension could not be ignored. The evaporation rate is suppressed by the weak interaction because the evaporation requires inverse-strangenization. We have only ever observed strong interacting matter with baryon number $A<300$ (atomic nuclei) or $A\approx 10^{57}$ (pulsars). If this assumption holds, the mass spectrum of strong interacting matter can be broadened considerably (Wu et al., 2024; Xu, 2024), see Figure 5.
@@
SnNs with $10^{25}<A<10^{35}$, as quark nuggets, are among the candidates for dark matter. Since the light-three-flavor symmetry in SnM is only slightly broken by the larger strange-quark mass, the charge-to-mass ratio of SnM is low, the electromagnetic signals from SnN are too weak to be observed. SnNs move virially in the universe at a typical speed of about $10^{-3}c$, when they hit Earth, they may not generate a high-energy shower signal, but rather an acoustic shock wave. So, we can detect these non-relativistic nuggets using the acoustic method (Qi and Xu, 2025). Due to the extremely high density of SnNs, their interaction cross-section with the ordinary matter is small. The cross-section of strangeon nuggets is defined as $\sigma_{\rm SN}=k~\cdot~(A/10^{30})^{2/3}\cdot 8.0\times 10^{-6}\,{\rm cm^{2}}$, in which $(A/10^{30})^{2/3}\cdot 8.0\times 10^{-6}\,{\rm cm^{2}}$ is the geometric cross-section. Therefore, SnNs’ penetration depth in rock or water is very large, even exceeding the radius of the Earth. The calculation results are shown in Figure 6.
@@
The initial shock wave is strong and sharp but will decay and widen during propagation in the medium due to diffusion, dispersion, dissipation and scattering. The decay can be described as an exponential function, and the frequency drop can be described as a polytropic function of the propagation distance, such as $R^{-1/4}$ or $R^{-1}$, depending on the medium. At a distance about 1 km, the frequency of the pulse signal is about 0.5 kHz to 5 MHz. At a critical distance $R_{\rm crit}$, the acoustic signal decay to the level of surrounding noise, $R_{\rm crit}$ is the maximum detectable distance. This distance depends on the state, the temperature and the uniformity of the medium. An estimated result is in Figure 7, $R_{\rm crit}\simeq 30\,$km. More accurate results require on-site acoustic experiments.